The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 3.00 $7.00 $21.00

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 250 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 1,272 F
Labor spending variance $ 2,812 F

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 2812-1272 = 1540 F

a) Labor efficiency variance = (Standard hour-actual hour)Standard rate

1540 = (250*3*7-7X)

7X = 5250-1540

X(actual hour) = 530 Hour

b) Labor rate variance = (Standard rate-actual rate)actual hour

1272 = (7*530-530X)

530X = 3710-1272

x(actual rate) = 4.6 per hour